OpenStudy (anonymous):

The no of possible st line passing through (2,3) and forming a triangle with coordinates axis, whose area is 12 sq. Units is a) 1 b) 2 c) 3 d) 4

OpenStudy (anonymous):

Draw it...

OpenStudy (anonymous):

What ?

OpenStudy (anonymous):

Let it pass through (a,0),(0,b) and (2,3).Frame an equation with these three points that also satisfy ab/2=12

OpenStudy (anonymous):

How beezlebub please explain the whole answer is

OpenStudy (anonymous):

the eqaution in intercept form is x/a + y/b=1 then even (2,3) fits this equation so 2/a + 3/b=1 and you have the second equation connecting the area, which is, ab/2=12 So now you have 2 eqautions to unknowns.Solve for a and b.

OpenStudy (anonymous):

2 equations 2 unknowns*

OpenStudy (anonymous):

ok let me solve it

OpenStudy (anonymous):

but how will i solve them

OpenStudy (anonymous):

Will u please help me

OpenStudy (anonymous):

a=24/b Now put that in the first equation, 2/(24/b)+3/b=1 b/12+3/b=1 Can you solve from here?

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

b = 6 a = 4

OpenStudy (anonymous):

Am i right ??????

OpenStudy (anonymous):

or b=4 and a=6. The equation is a quadratic so you should get 2 answers fro b, and likewise two for a

OpenStudy (anonymous):

b=4,a=6 or b=6,a=4

OpenStudy (anonymous):

Right Now what will I have to do ?

OpenStudy (anonymous):

so now depending on the values or a and b you plug in, you get 2 equations.So 2 possible straight lines.

OpenStudy (anonymous):

beezlebub thanks a lot for help Will u help me in 1 more question ? Please

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

Go to this link http://openstudy.com/groups/mathematics/updates/4e3169b50b8ba7b2da40fe0d

OpenStudy (anonymous):

Beezlebub are u solving the question ?